Zippy Shoe Co. uses a periodic inventory system. Zippy purchased 405 pairs of shoes at $68 each in June, 970 pairs in August at $70 each, and 620 pairs in December at $73 each. Zippy sold 1,890 pairs of shoes during the year.Zippy Shoe Co. uses a periodic inventory system. Zippy purchased 405 pairs of shoes at $68 each in June, 970 pairs in AugustComplete this question by entering your answers in the tabs below. Required A Required B Required LIFO Cost of Goods Sold EndComplete this question by entering your answers in the tabs below. Required A Required B Required C Weighted Average (Round y
Answer #1
Unit Unit Cost Total Cost
Purchase 1 405 68 27540
Purchase 2 970 70 67900
Purchase 3 620 73 45260
1995 140700

a) Calculate following : FIFO

Cost of goods sold = 27540+67900+515*73 = 133035

Ending inventory = 105*73 = 7665

b) Calculate following : LIFO

Cost of goods sold = 45260+67900+300*68 = 133560

Ending inventory = 105*68 = 7140

c) Calculate following: Average cost

Cost of goods sold = 70.53*1890 = 133302

Ending inventory = 140700-133302 = 7398